construct a quadratic equation whose solutions are -5, 7
Answer by Edwin McCravy(20055)

I'll give you the equation first, then I'll tell you how I got it.
The equation is x²-2x-35 = 0.

If we were asked to solve that, we would do this:

                    x²-2x-35 = 0
                  (x+5)(x-7) = 0
                x+5 = 0;   x-7 = 0
                  x = -5;    x = 7


Just reverse those steps and you'll get how to go from the solutions
back to the equation.  You are "solving it in reverse":

                  x = -5;    x = 7
                x+5 = 0;   x-7 = 0
                  (x+5)(x-7) = 0
                    x²-2x-35 = 0

Edwin